OpenStudy (anonymous):

I need some help, what is the integral of (e^.2t)/(4+e^.2t) Thank you in advance!

OpenStudy (amistre64):

the top is practically the derivative of the bottom ... what rule does that follow?

OpenStudy (amistre64):

or if a u sub is more up your alley ... let u = the denominator and work out the subsitutions

OpenStudy (anonymous):

so if you take the u subsitutation is the denominator the du is just the top. So the answer would be 1/e^.2t?

OpenStudy (amistre64):

let u = 4+e^.2t du = .2 e^.2t dt du/.2 = e^.2t dt \[\int\frac{1}{.2}\frac{du}{u}\]

OpenStudy (amistre64):

ln(4+e^.2t)/.2 + C is what im lookng at